La Silla Paranal Observatory News
- Nov 20, 2013: recoating of M1, M3 of UT2 completed
- Sept 11, 2013: recoating of M1, M3 of UT4 completed
- April 1, 2013: The VST is currently out of service between March 29, 2013, and April 2, 2013 due to a multiple damage of cables and cooling lines after an incident caused by the failure of a limit switch. All damaged parts were replaced or repaired.
- February 26, 2013: The Laser Guide Star system at UT4 has recently been successfully upgraded with a fibre Raman laser (AOF prototype) delivering improved on-sky performances as well as enhanced operational stability and maintenance. Full characterization of the upgraded system will soon be available. This intervention will result in a dramatic increased in the availability of the LGS system for science operations.
- September 19 to October 8, 2012: UT4/Yepun was out of operations due to a serious problem with the telescope hydraulics. The repair was completed on Oct. 7/8 and the telescope was recalibrated and handed over to operations in a very good shape.
- September 7, 2012: recoating of M1, M3 of UT3 completed
- June 11, 2012: recoating of M1, M3 of UT1 completed
- June 4, 2012: UT4 - Yepun: The UT4 Laser Guide Star (LGS) system will be unavailable for operations in January and February 2013. It will be replaced by a prototype laser for the Adaptive Optics Facility (AOF). This new laser system is expected to provide easier maintenance and operations, and the number of LGS nights available should significantly increase.
- March 21 - May 3: UT4 - Yepun: Because of preparatory work for the installation of the Adaptive Optics Facility, Yepun is not available for operations.
- November 17, 2011: recoating of M1, M3 of UT2 completed
- August 13, 2011: Start of VST Public Surveys with Early Science and Dry Runs in parallel to technical activities
- March 14 - May 4, 2011: Extended VISTA recovery period. A serious mechanical problem detected during the coating intervention requires an extended repair and preventive maintenance period. The VISTA telescope is expected to be back for on-sky tests by May 3/4, 2011.
- March 4 - April 7, 2011: Recoating of M1 and M2 of VISTA with aluminium.
- December, 2010:Commissioning of the VST started with all systems integrated. A smalltest CCD is installed in place of OMEGACAM (the science instrument).OMEGACAM will be installed at the VST in March 2011.
- November 23, 2010: recoating of M1 of UT1 completed
- September 28, 2010: recoating of M1 of UT3 completed
- November 3, 2009: recoating of M1 of UT2 completed
- September 8, 2009: recoating of M1 of UT4 completed
- June, 2009:
The last part of the VST, the M1 cell, arrived to Paranal this month. Engineering activities will now continuously ramp up throughout the second semester of 2009, aiming for a start of operations of the VST in 2010.
Photo G. Hüdepohl
HAWK-I on NAS B of UT4 ready for its first commissioning
- November 13, 2008: recoating of M1 of UT1 completed
- November, 2008: After a successful period of technical activities, the LGS facility is now fully offered with both NACO and SINFONI.
- August, 28, 2008: Following the LGS recovery actions which took place at the end of July, it has been decided to resume operations of SINFONI with the laser (service mode only). The use of the laser with NACO is still waiting for the results of a commissioning run scheduled for November 2008.Moreover, due to the intrinsic delays in getting the LGS system in operations, the commissioning of the seeing-enhancer mode with SINFONI has been postponed to early 2009. Updated news will be posted on this webpage when they become available.
- June, 26, 2008: Although substantial improvements were made to the LGS during the recent technical activities that took place at UT4, it was decided that an additional set of recovery actions was needed - to be scheduled end of July 2008. Until then, the LGSF remains unavailable for science operations.Please, continue to check this webpage for the latest news on the LGSF recovery activities.
- May, 23, 2008: LGSF recovery activities have been extended until mid-June 2008.As a result, the LGSF remains unavailable for scienceoperations. Depending on the success of the June LGS recovery,executions of some science programs in Service Mode couldresume by the end of July 2008.
News on the ongoing recovery activities will continue to be posted here.
- April, 30, 2008: The continued degradation of the performance, reliability, and availability of the Laser Guide Star Facility (LGSF) has forced to Observatory to take this facility out of operation as of April 1st, 2008. All service mode programs requiring the LGS are on hold. Visitor runs with the LGS will have to be cancelled.
After a deep analysis of the situation a detailed and comprehensive recovery plan has been developed by ESO together with the LGS partner institutes to remedy this situation. The plan was released on April 25 and foresees the implementation of immediate short-term actions for the period April - July to recover the LGSF performance as required for successful science operation (i.e., the performance as of June 2007).
A first critical milestone of the recovery plan is to have the LGSF operational for mid-May where a LGS commissioning together with NACO has been scheduled earlier.
In parallel to the implementation of the short-term plan a medium-term plan for the LGSF is developed to improve its reliability and performance on system level. This plan is expected to be released in October this year.
- April 1st, 2008: AMBER: Fringe tracking with FINITO is availabe to visitor mode PIs
- April 1st, 2008: HAWK-I: start of Science Operation
- July 2, 2007: recoating of M1 of UT1 successfully completed
- March 8, 2007: recoating of M2 of UT2 successfully completed
- January 2007: AT4 commissioned and accepted for operations
- December 5, 2006: UT2 recoating of M1/M3 completed
- September 5, 2006: UT4 recoating of M1/M3 completed
- May 12, 2006: First recoating of M2 of UT1 successfully completed
- August 24: KMOS arm 7 is parked, locked and in simulation
- August 6, 7: KMOS arm 5 is parked, locked and in simulation; arm 15 is recovered and back in operation
- July 17: last night of operation for CRIRES before the start of the upgrade
- July 7: on July 4, an anomalous noise was heard coming from inside the OMEGACAM's cryostat. By precaution, the instrument was out of operation on the nights of July 4 and 5th.
- June 20 - 29: MUSE Science Verification
- June 4: press release describing the first light of SPHERE.
- June 2: speeds of the grating wheel of SINFONI had to be decreased
- May 31: KMOS arm 15 is parked, locked and in simulation
- May 20: KMOS arms 4 and 11 are temporarily out of operations.
- May 20: FLAMES ETC with improved efficiency measures at all settings.
- May 10: start of the SPHERE commissioning.
- April 1: KMOS Arm 11 is back in operation.
- March 5: press release describing the first light of MUSE.
- March 1: Limiting magnitudes for AMBER in the K-band has increased, owing to polarization control.
- February 21: KMOS was successfully recommissioned after an intervention to lower the vibration and to repair the arms that were locked during the last months. Arm 11 is still not available, all other arms work well within the specifications.
- February 7: start of the MUSE commissioning.
- January 7: KMOS arm 11 is parked, locked and in simulation.
- January 7: UVES red arm CCD shutter failure. See the UVES News page.
- December 23: KMOS arm 10 is parked, locked and in simulation
- October 15: KMOS arm 2 is parked, locked and in simulation
- October 1: KMOS is now in operation
- HAWKI on UT4 is now out of service, due to an intervention aimed at cleaning the entrance window. The intervention will be completed by next October 23, 2013.
- March 11: SINFONI: The motors for the Grating and Filter wheels are failing. An intervention is planned for September 2013. In the meantime, users are asked to reduce the number of spectrograph configurations as much as possible. See SINFONI news for details.
- Feburary 27: VISIR is not offered neither in P91 nor in P92, as a better understanding of the detector issues is needed. See the VISIR Upgrade Project page.
- January 17: Start of the 2nd KMOS commissioning.
- December 18: FLAMES-UVES: Fiber 37, plate 2 broke on September 11. Latest attempt to fix it yielded 30% of the original transmission. Therefore the fiber will remain disabled until better results are reached or the fiber is replaced. Follow-up on FLAMES news
- December 4: The first commissioning of KMOS has been successfully completed.
- November 20: NACO intervention is completed. See NACO news for details.
- September 7, 2012: NACO is foreseen to be refurbished from October 1, to mid-November 2012
- September 7, 2012: NACO: Between August 25th and September 5th, issues with the NACO field selector electronics prevented the use of the IR WFS. More details in the NACO news.
- August 16: VISIR: The first commissioning of the upgraded VISIR took place between July 28 and August 10. Results are mixed. The prism provides the expected coverage and resolution. Good progress has been made on the characterisation of the new modes. Unfortunately, the imager does not provide the expected throughput. The instrument will be brought back to the integration hall before the M1 coating (Aug 27 - Sep 6). A search for the cause of the decreased sensitivity will be carried out in addition to planned modifications. It is foreseen that the instrument will not be operational before end of October. More details in the VISIR Upgrade Project page.
- August 16: FORS2: The 'CHIP 1' (or upper chip) of the Red CCD mosaic failed in the morning of August 8, 2012. The Red CCD mosaic was replaced by the Blue one on August 10 to allow for a better set-up diagnostics on the CCD test bench. Several problems were found and solved. The Red CCD mosaic was put back on the instrument on August 16.
- August 16: XSHOOTER: Recently the rate of XSHOOTER ADC failures have increased. There is an ongoing investigation to find the cause for this misbehaviour, but it is unlikely that the situation is back to normal for the next few months. A major intervention on XSHOOTER to fix this problem is under evaluation. Incorrect position of ADCs might lead to slit losses worse than if they are not used. Consequently, the ADCs were temporarily disabled, starting on August 1st, 2012. More details in the XSHOOTER News page.
- June 4: KMOS, the VLT 2nd Generation K-band multi-object spectrograph, will start to be commissioned in November 2012.
- March 21 - May 3: UT4 - Yepun: Because of preparatory work for the installation of the Adaptive Optics Facility, Yepun is not available for operations. Observations with HAWK-I, NACO and SINFONI are therefore not possible.
- February 15, 2012: NACO: the detector of CONICA suffers strong electronic patterns affecting the two upper quadrants. See the NACO News page for details.
- October 15, 2011: OmegaCAM enters in operations.
- October 13, 2011: CRIRES is back in operations following an internvention whose main goal was to replace the entrance slit mechanism of CRIRES. CRIRES has now 2 fixed slits, of widths corresponding to 0.2 and 0.4 arcsec on sky, engraved in a gold coated silicon wafer plate by microlithography and etching.
- July 12, 2011: ISAAC is back: a small piece of metal had managed its way towards the connectors of the Hawaii detector and caused ashort circuit. It could be removed and fortunately the detector was not damaged.
- June 3, 2011: Since May 18, ISAAC can only be operated using the long-wavelength Aladdindetector. The spare Hawaii detector is being tested and should be installed in ISAAC soon. ISAAC should be warmed up on June 6. The instrument should be back in operations by the end of June.
- May 18, 2011: The Hawaii detector used in the ShortWavelength arm of ISAAC has 'lost' one quadrant. The suspected natureof the fault suggests that it should be switched off for safety reason.
- April 28, 2011: The intervention and recommissioning of the NACO visible wavefront sensor were successfull. The Visible WFS is nowback in operations.
- March, 2011: Start of OmegaCAM commissioning.
- March 20: A high-risk earthquake hit Paranal affecting UVES(image slicer alignment), ISAAC (rotator offset changed by 0.1deg).
- February, March. The selector of the visible wavefrontsensor of NACO is unreliable. As of March 6, only the visible 14x14 wavefrontsensor is available, preventing in particular the use of NACO with theLaser Guide Star Facility. An intervention is planned to restore thefullchoice by April 24. The IR WFS is not affected.
- November 22-25: UVES: The turntable for the grating of the UVES blue-arm has been replaced.
- August, September: VIMOS is back on sky, but with problems. See the VIMOS News page.
- May 22 - July 31: VIMOS is taken out of operations forupgrade, in particular, with new red-sensitive detectors and AutomaticFlexure Compensation.
- July 16: CRIRES: A new intervention improved the parallelism of the entrance slit and its alignment with the intermediate slit.
- June: MIDI: The readout windows were shifted by 7 pixels, to recenter the PRISM spectra.
- May 01 until July 10: CRIRES: Problems affecting the coolingmade the instrument unavailable for most of May during which a newcompressor was installed. Unfortunately, following the intervention,the entrance slit lost its parallelism and its alignment with theintermediate slit.
- May: SINFONI: A new STRAP, without central obscuration, was installed. Commissioning hampered by bad weather.
- April 2-3: NACO: The Apodized Phase Plate coronograph was successfully commissioned.
- March 30: CRIRES: The MACAO shutter was replaced.
- Feburary 20: FLAMES-GIRAFFE: Maintenance and repair of CCD shutter.
- Feburary 13-15: ISAAC: Filter wheel 2 was repaired.
- Feburary 7: NACO: Successfull installation and commissioning of a 14x14 Visible WFS for use with the Lase Guide Star Facility.
- October, 2009: XSHOOTER has been fully commissioned and installed at the Cassegrain of UT2; the instrumentofficially started scientific operations from October 1st.
- September, 2009: ISAAC is back in operations after to be re-installed at the Nasmith A of UT3.
- August, 2009: ISAAC has been succesfully installed at the Nasmith A of UT3; the recommissioning on sky is currently finalizing.The Instrument should be back in operations before the end of the month.
- July, 2009: The UVES red-arm has succesfully recommissioned and the instrument is fully back in operations fromJuly 22nd. The UVES red-arm is scheduled to be recomissioned July 10-12 and will be offered onward. ISAAC will be removed from UT1 at the end of this month, and remounted to UT3 early September to help balancing pressure factor across the UTs.
- April, 2009: The two FORSes have ben merged into a single instrument, FORS2,which is still offered at UT1, but with polarimetric capabilities,starting P83. X-SHOOTER has been moved to UT2 for continuation of commissioning and is scheduled to start operations next semester (P84). VISIR has been reinstalled at the UT3 Cassegrain focus and is fully back in operations.
- October, 2008: HAWKI has been fully commissioned and installed at the Nasmith focus of UT4; the instrument officially startedscientific operations from October 1st.
- August 1st, 2007: HAWKI: First Light of the instrument mounted at UT4.
- April 1st, 2007: AMBER: start of Science Operation with 3 ATs. Fringe tracking with FINITO is availabe to visitor mode PIs
- April 1st, 2007: CRIRES: start of Science Operation
- April 7, 2007: FORS1 CCD upgrade completed
- March 25, 2007: First light for MAD, the multi-conjugate adaptive optics demonstrator
- March 13, 2007: SINFONI: first commissioning with LGS completed
- March 8, 2007: Installation of new FLAMES FACB guiding TCCDs completed
- January 2007: NACO: first commissioning with LGS completed
- December 18, 2006: LGSF successfully commissioned
- December 5, 2006: Installation of new UVES slit viewers TCCDs completed
- October 31, 2006: DAZLE visitor instrument succesfully commissioned on UT3
- October 29, 2006: Launch Telescope of LGSF successfully re-installed
- June 2006: commissioning of CRIRES at UT1 started
- April 2006: CRIRES AO system (MACAO) installed and commissioned at UT1
- January 28, 2006: first light of the Laser Guide Star Facility LGSF at UT4
On December 16 at 05:10 LT another earthquake with magnitude 6.8 ( Richter) affected the north of Chile. In Paranal it was registered with a magnitude of 5.2 and a peak ground acceleration of 0.2m/s2. Neither injuries to the people nor damages to the installations have been registered.
On November 14 at 12:41 LT an earthquake with magnitude 7.7 (Richter) with epicenter north of Antofagasta affected the north of Chile. In Paranal it was registered with a magnitude of 5.7, a peak ground acceleration of 0.6m/s2 and for a duration of approx 2 minutes. This has been the strongest earthquake since 1997, the first "high risk" earthquake since the start of operation of the VLT, and the first strong earthquake since the opening of the residencia. Neither injuries to the people nor damages to the installations have been registered. After extended checkout procedures, the regular night time operations have been resumed at the sunset. All the systems have been operated without restrictions, including all four UTs (with their instruments and the LGS) and the ATs with VLTI.
Several other minor episodes have been registered in the following 24 hours without any impact on the normal VLT activities.
The section provides the latest telescope and instruments news from the La Silla Site.
More news can be found in the news section of theLa Silla Science Operations page.
- March 14, 2010: Updated service observing information Service Observing
- November 20, 2009: A new technique using near-infrared images, obtained with s 3.58-metre New Technology Telescope (NTT)ESO Watching a Cannibal Galaxy Dine
- April 6, 2009: La Silla Astroclimatology information (weather, seeing, extinction) was moved to a new location and can now be found on the mean navigation menu of the La Silla Facilities pages. Additional La Silla astroclimatology information will be added here.
- February 11, 2009: From the second week of July, 2009, observations will be carried out from a new control room located in theadministration building.
- February 11, 2009: A new page describing the way that operations will run after P83 (ie from October 2009) is available here.
- April 22, 2008: A document that explains how to interface visitor instruments with focal stations of La Silla telescopes is available (2.8 MB PDF format).
- March 27, 2007: The dome of the 3.6m has been fully restored, and moves with nominal speed.
- October 05, 2006: The dome of the 3.6m enters an extended period of corrective maintenance. Science operations at the 3.6m is temporarily suspended until further notice.
- May 17, 2006: A Trio of Neptunes discovered with HARPS, see ESO PR 18/06
- January 26, 2006: A 5.5M_earth planet discovered with the 1.54m Danish telescope.